OpenStudy (anonymous):

Find the value of a in the equation 5 / a + 3 = 3 / a − 2.

OpenStudy (anonymous):

is a+3 and a-2 are in denominator??

OpenStudy (mertsj):

Cross multiply

OpenStudy (anonymous):

they are in denominators

OpenStudy (mertsj):

Cross multiply

OpenStudy (anonymous):

yeah,,,mertsj is right....

OpenStudy (anonymous):

im not understanding how to do it

OpenStudy (mertsj):

\[\frac{5}{a+3}=\frac{3}{a-2}\]

OpenStudy (mertsj):

\[5(a-2)=3(a+3)\]

OpenStudy (anonymous):

 you cross multiplt, then simplify. . .